OpenStudy (anonymous):

Find the instantaneous rate of change of w with respect to z if w=7/3z^2.

zepdrix (zepdrix):

They want us to find \[\large \frac{dw}{dz}\] So we need to take the derivative with respect to z. Confused about how to do that? :o Have you learned about derivatives yet?

OpenStudy (anonymous):

Well I know how to take derivatives, I just wasn't sure about how to go about it. Would I use the... well, I think it's called the quantitative rule but I'm not sure. It's ba'-ab'/b^2. So that would be ((3z^2*0)-(7)(6zz'))/3z^4, right?

OpenStudy (anonymous):

Without simplifying, of course.

zepdrix (zepdrix):

The Quotient rule? You certainly could, but that seems unnecessary since the top term is a constant. It's probably a little bit better of an idea to - rewrite the z term with a negative exponent. Remember how to do that? You bring it up to the top from the denominator, and change the sign on the exponent.

zepdrix (zepdrix):

\[\huge \frac{7}{3z^2}\qquad \rightarrow \qquad \frac{7z^{-2}}{3}\]Which can written even nicer like this,\[\huge \frac{7}{3}z^{-2}\]From here it's just the power rule.

zepdrix (zepdrix):

The way you did it looks good though :) I don't see any mistakes. So if that method makes more sense to you, then by all means stick with it. Oh just remember that when you SQUARE the b term in the bottom, you square the entire b, which would include the 3.

zepdrix (zepdrix):

So maybe one lil mistake heh.

OpenStudy (anonymous):

Ahh okay, I didn't even think of it that way! I forget about all of my options sometimes, haha. So, simplified it would be\[w'=\frac{ -14z' }{ 3z^3 }\] Right?

zepdrix (zepdrix):

Yah looks good c:

OpenStudy (anonymous):

Okay, thanks so much! :)